Hi guys, Firstly I am new to this forum but it is amazing. I am an 11th grade student doing PCM and want to clear JEE-advanced. I have joined an Institute called Wisdom here. I am nothing more than an average student. Can anyone make a daily routine for me which I can follow for studies? My school is 8am to 12:30 pm and my tuition is on Monday, Wednesday, friday and satyrday from 2pm to 7pm. Thanks in Advance.

Grade:11

1 Answers

Umakant biswal
7 years ago
@ paras 
thanks for being a part of the forum . 
as u are a grade 11th student and want to clear jee advanced , u have ample of time to do so , and u have also a reduced school hours , normally its 6 hrs for major people . 
so, u can use that for study . 
TIME TABLE TO FOLLOW 
6 AM – WAKE UP 
6- 7 AM – YOGA / FRESH UP 
7 -7;45 – READY FOR SCHOOL 
8 – 12 = FOCUS WHATEVER TAUGHT ON SCHOOL 
12;10- 1;30 – DAYTIME MEALS . 
1;45 – 3;00 – TAKE REST 
3;00-3;15 – FRESH UP 
3;20 – 4;30 = physics ( revision of concept )
4;;30 – 5;30 – doing nuericals of the studied chap[ter ( phy ) 
5;30 – 5 ;45 – break 
5;45 – 7 ; 45 = maths ( solve at moist ques as u can ) 
7;45 – 8 ;00 – break 
8;05 -9;30 – chemistry (in your tution day do inorganic and organ ic , and non tution day do physical) 
9;30-10;00 – english 
10;00 – 11;00 – tv watching / night meal . 
11;05 – 12;05 – recap phy and maths formula that u have studied . 
12;10 – 12;30 – solve previous year ques paper ( jee main / advanced)) 
12; 35 –  5;55 – sleep and wake up 
THIS IS THE TIME TABLE FOR NON TUTION DAYS , IN TUTION DAYS MAKE NESSASRY ARRENGEMENT WITH THE SUBJECT THAT U LEARNT IN TUTION , 
AND make sure give max time to weak subject , u can reframe the time table if u find less time for weak subject . 
ALL THE BEST ..